# LoopSR: Looping Sim-and-Real for Lifelong Policy Adaptation of Legged Robots

### v1

@@ Abstract
Reinforcement Learning (RL) has shown its remarkable and generalizable capability in legged locomotion through sim-to-real transfer. However, while adaptive methods like domain randomization are expected to make policy more robust to diverse environments, such comprehensiveness potentially detracts from the policy’s performance in any specific environment according to the No Free Lunch theorem, leading to a suboptimal solution once deployed in the real world. To address this issue, we propose a lifelong policy adaptation framework named LoopSR, which utilizes a transformer-based encoder to project real-world trajectories into a latent space, and accordingly reconstruct the real-world environments back in simulation for further improvement. Autoencoder architecture and contrastive learning methods are adopted to better extract the characteristics of real-world dynamics. The simulation parameters for continual training are derived by combining predicted parameters from the decoder with retrieved parameters from the simulation trajectory dataset. By leveraging the continual training, LoopSR achieves superior data efficiency compared with strong baselines, with only a limited amount of data to yield eminent performance in both sim-to-sim and sim-to-real experiments.
@@ I Introduction
Reinforcement Learning (RL) has seen significant progress as a viable alternative for robotic control, but applying it to real-world scenarios remains challenging. Advances in simulation systems [1, 2] and extensive parallel training [3] have enhanced policy performance and robustness, but the sim-to-real gap persists. Adaptive methods such as domain randomization [4, 5, 6, 7, 8] are widely used to address the gap and improve robustness. However, given RL methods’ inherent requirement of direct interaction with the environment, and along with No Free Lunch Theorem [9, 10], which suggests a potential trade-off between generalization and specific performance, purely training in the simulation with inaccurate state occupancy distribution will limit the policy’s performance beforehand.
@@
Leveraging the data in the real world where the robot is directly situated is an intuitive solution to such a defect. Nevertheless, it still faces obstacles. Firstly, collecting real-world data is notoriously expensive while RL methods are extremely data-hungry. An applicable policy generally requires several months of real-world experience which is hardly affordable. Secondly, privileged knowledge (i.e. heightfield, contact force, mass, friction, etc) is absent in real-world settings, causing troubles when facing tricky terrains. One example is the stairs, where a blind robot with only proprioceptive observations (i.e. joint position & velocity) needs extensive exploration to learn to lift its legs. while the one with precise height information can better find such a solution. On top of that, noisy observation of the real world leads to training instability.
@@
Previous works have attempted to cope with these impediments by reshaping reward functions [11, 12], taking advantage of sample-efficient algorithms [11, 12, 13], or harnessing model-based RL [14]. However, these existing methods, trained directly in the real world, fail to generate superior locomotion policies with safe and steady performance compared with prevailing zero-shot methods and are vulnerable in training.
@@
Drawing intuition from the cognitive process where animals tend to leverage the experience in familiar environments when encountering new ones, we investigate to find a bridge back to the controllable simulation environment instead of trapping ourselves in the capricious and noisy real world. The benefits of such an idea are all-round. In one way, we can sidestep the bothering problem of redesigning a delicate reward function corresponding to the limited observation, and take full advantage of a well-polished simulation training process [1, 3, 4, 15, 16]. Additionally, with such a method, we only require a minimal amount of data to identify the surroundings well and still live off the benefits of data abundance in the simulation.
@@
In this work, we present LoopSR, an effective pipeline for lifelong policy adaptation of legged robots depicted in Fig. 1. The core of LoopSR is a transformer-based encoder [17] to leverage the latent space. Particularly, we adopt autoencoder architecture [18, 19] and contrastive loss [20, 21, 22, 23, 24, 25] to guide the model to extract sufficient features pertinent to reconstructing the simulation, ensuring the effectiveness and generalizability of the latent space. We implement both learning-based and retrieval-based methods to derive simulation parameters from the latent variable, which compensate for each other in accuracy and robustness to out-of-distribution scenes. LoopSR is tested in both IsaacGym simulated environment [1] and the real-world scene compared with strong baselines.
@@
We proposed a novel lifelong policy adaptation method LoopSR that first incorporates real-world continual learning into simulated policy pretraining, achieving constant improvement during real-world deployments.
@@ II-A Representation Learning in RL
Representation learning [19, 26, 27, 28] has made up the bedrock for modern deep learning algorithms, and the field of RL and control is no exception. Previous works have made versatile attempts to take advantage of the representation learning and latent space [29, 30, 31], including contrastive learning [25, 32], reward function encoding [33, 34], domain/task-specific representation [18, 35, 36, 37], and gait encodings [38, 39], boosting multi-modal input [27] or transferring through different tasks and embodiment [36].
@@ II-B Learning in the Real World
Continual learning in RL [42] is a daunting task as the result of the compounding error in drifted distribution. Works have been done to debase the trouble caused by protean data distribution through experience replay [43, 44], distillation [45], and reusing parts of model [46], etc. However, when it comes to real-world learning, the situation becomes progressively difficult due to noisy input, limited observation, and inaccessible reward functions.
[11, 12, 13] are among the few works tapping into this tricky problem. Instead of sim-to-real transfer, Smith et al. have decided to completely learn the policy in the real world, and have made attempts to introduce reshaped reward functions and techniques of restricting action space to boost effective training. However, these algorithm needs careful designing in action space constraints and rewards to reluctantly obtain effective training and executable gaits, while failing to outstrip zero-shot methods.
With world models exhibiting competence in heterogeneous downstream tasks [47, 48, 49], [14] employ the Dreamer world model in the real-world learning task. The model-based method, though a practical alternative, still suffers from the problem of deficient privileged knowledge. The model’s accuracy in featuring real-world information is also worth questioning in comparison with the simulation, as model-based methods are susceptible to pervasive model biases. The vulnerability of training also places the risk that convergence is not guaranteed facing the composition of model biases and compounding error.
@@ III Methodology
III-A Priliminaries
We formulate the training process in the simulator as a Partially Observable Markov Decision Process (POMDP) as the environment cannot be fully observed. We define the MDP in the simulation as , where denotes the state space, observation space, and action space respectively. denotes the transition function, and the reward function is denoted as . denotes the discount factor. We define the privileged simulation parameters used to construct the simulation environment. Specifically, contains the knowledge of the terrain distribution (terrains are divided into 5 categories for feasibility) and robot-concerned parameters (i.e. mass, restitution, motor strength). The goal of the RL method is to get the policy to maximize the expected discounted return:
In the task of the lifelong policy adaptation, we need to find the optimal policy with the pretrained one in the real scene with the POMDP , where is implicit without access. In the real-world setting, we are able to constantly collect trajectories but at an extremely low speed at least times slower than in the simulation.
In the later analysis, we take and to emblematize the training environment and the testing one.
@@ III-B Looping Sim and Real
We present our system for quadrupedal lifelong policy adaptation as Looping Sim and Real (LoopSR). The overview of LoopSR is depicted in Fig. 1. We draw on Proximal Policy Optimization (PPO) [50] to pretrain the policy in the IsaacGym environment following [3]. The policy is constructed following the state-of-the-art algorithm for quadrupedal locomotion [51, 52, 53], utilizing an Asymmetric Actor-Critic (AAC) architecture and expert motion data to enable robust sim-to-real transfer. Domain randomizations [6, 7, 8] are added to introduce a robust policy.
We used the collected trajectories with the environment parameters (privileged information & initial domain randomization range) during the training process to learn LoopSR-enc which projects the whole trajectory to the latent space for relevant information. A multi-head decoder is designed to estimate the privileged information from the latent variable. In testing environments where the robot is deployed, the collected trajectory is encoded as the latent variable . The estimated is the weighted sum of the retrieval results referring to the training dataset of LoopSR-enc and the decoded results , which is returned to IsaacGym environment to establish the simulation accordingly. Some techniques are taken to maintain stable training illustrated in Sec. III-D. The pseudocode 1 summarizes the whole pipeline.
@@ III-C Network Architecture
Contrastive Loss. To guide the model to implicitly discriminate different terrains, we apply the contrastive loss with the form of supervised InfoNCE loss similar to [23]. The loss is defined to maximize the inner product of from the trajectories with the same terrain type label, whose theoretical background of maximizing the mutual information between positive samples and the latent representation is well-discussed in [22, 23].
where is the batch size.
Multi-head Decoder. We also design a multi-head decoder and to separately extract the terrain distribution and the robot proprioceptive parameters for disentanglement of different environment characteristics. We introduce , , and to represent the environment terrain distribution, robot parameter, and the range of the domain randomization in the training process. The losses are designed as follows:
In Sec. IV-D, we conduct the ablation study to discuss the necessity of all loss terms.
@@ III-D Implementation Details
Sampling-based Policy decomposition. Rather than collecting training data exclusively with the last pretrained policy, the dataset comprises trajectories sampled from different policy checkpoints. In this way, a sampling-based policy decomposition is inferred, ensuring stability when trajectories are collected from a constantly updating policy.
Soft Update. We implement the soft update of estimated environment parameters to avoid deleterious abrupt changes in terrain distribution. The terrain distribution is updated through . is set to be . For the robot parameters, we half the domain randomization range to balance robustness and performance. Such a method has the adaptive ability to attune to the environment setting which is not in the initial reference dataset. The necessity of Soft update is illustrated in Sec. IV-D.
Retrieval Process. Regarding training safety, we use the retrieval method widely adopted in recommender systems [59] and the application of Large Language Models [60] to preclude drastically mistaken results in the face of out-of-distribution (o.o.d.) instances from learning-based methods. The final result is weighted as with in this work set to be .
@@ III-E Theoretical Analysis
We provide a theoretical analysis here to explain the effectiveness of LoopSR. With the No Free Lunch Theorem [9, 10] telling us a specific model may outstrip a general one, we demonstrate that the wide DR in policy training leads to suboptimal strategy in a specific environment. The discrepancy between optimal value functions in the training environment and the testing one is derived in Eq. 5.
where and are optimal value functions in and respectively, and as the optimal state value function on . informs the discrepancy in reward functions, models the discrepancy caused by policy difference with detailed reasoning in [61], and denotes the discrepancy induced by the shift in state occupancy distribution. The discrepancies are detailed as follows:
In LoopSR, is negligible as is the sound approximate for given the substantiated validity of [3, 4, 15, 16]. For the second term, according to [61], will approach with sufficient model capacity and iteration steps as and share the same input. For the last term, is closely related to the innate state occupancy distribution (e.g. the state distribution of the stair terrain differs from the plain one). Parameter can efficiently model such an error and mitigate it. To sum up, LoopSR can eliminate the discrepancy in Eq. 5 through effective modeling of each term, thus optimizing the policy performance.
@@ IV Experiments
| Terrain | Difficulty | Expert | Origin | LoopSR(ours) | PPO | SAC | World Model |
@@ IV-A Sim-to-Sim Experiment
Overview. Although simulation settings cannot perfectly replicate real-world complexities, they offer a controlled environment for conducting statistical experiments, facilitating comparisons when direct real-world computation is impractical. We leverage IsaacGym [1] for training and simulation, which provides implementations of various terrains. The training code is developed based on [3].
The experiment consists of two primary phases. The first phase is pretraining in the training environment, where 4,096 robots are simulated across five distinct terrains (upward/downward slopes/stairs and flat ground), training in parallel for 30,000 iterations, with 24 simulation steps per iteration. The second phase involves data collection and continual training in a testing environment chosen from three terrains, each with three difficulty levels, defined by parameters of inclination, step height, and obstacle height, respectively (see Table I). Privileged knowledge is not accessed in the testing environment to simulate real-world conditions better. Each policy is trained on steps of collected data, equivalent to half an hour in a real-world experiment. During evaluations, robots are commanded to walk forward at a velocity of to assess their maximum performance. The reward functions follow [3] and are averaged across simulation steps. LoopSR takes out sim-to-real loops for each terrain during simulation experiments.
@@
Expert is a reference policy completely trained on the desired test environment.
Origin is a reference policy without any refinement, representing a zero-shot sim-to-real workflow. It is trained for the same number of timesteps as the combined pretraining and continual training steps.
PPO is the vanilla PPO method without the privileged knowledge with a reward function the same as Origin, excluding privileged terms.
SAC follows [11] with SAC used to pretrain and continual training. We modify the target velocity in the original reward function to for testing.
Results. The simulation results are shown in Tab. I, where LoopSR achieves prominent enhancement and closely matches the performance of Expert, especially in challenging terrains like Stairs and Discrete. This outperformance over Origin confirms LoopSR’s success in continuous optimization. At the same time, direct real-world training approaches (PPO, SAC, World Model) mainly fail such an aim, typically owning to their fragility to the shift in state occupancy distribution. The reward curve concerning the sim-to-real loops is shown in Fig. 3, where each sim-to-real loop integrates simulated training of iterations. We find out that LoopSR surpasses the performance bottleneck which troubles the original baseline.

### v2

@@ Abstract
Reinforcement Learning (RL) has shown its remarkable and generalizable capability in legged locomotion through sim-to-real transfer. However, while adaptive methods like domain randomization are expected to enhance policy robustness across diverse environments, they potentially compromise the policy’s performance in any specific environment, leading to suboptimal real-world deployment due to the No Free Lunch theorem. To address this, we propose LoopSR, a lifelong policy adaptation framework that continuously refines RL policies in the post-deployment stage. LoopSR employs a transformer-based encoder to map real-world trajectories into a latent space and reconstruct a digital twin of the real world for further improvement. Autoencoder architecture and contrastive learning methods are adopted to enhance feature extraction of real-world dynamics. Simulation parameters for continual training are derived by combining predicted values from the decoder with retrieved parameters from a pre-collected simulation trajectory dataset. By leveraging simulated continual training, LoopSR achieves superior data efficiency compared with strong baselines, yielding eminent performance with limited data in both sim-to-sim and sim-to-real experiments.
@@ I Introduction
Reinforcement Learning (RL) has seen significant progress as a viable alternative for robotic control, but applying it to real-world scenarios remains challenging. Advances in simulation systems [1, 2] and extensive parallel training [3] have enhanced policy performance and robustness, but the sim-to-real gap persists.
Adaptive methods such as domain randomization [4, 5, 6, 7, 8] are widely used to address the gap and improve robustness. However, given RL methods’ inherent requirement of direct interaction with the environment, and along with No Free Lunch Theorem [9, 10], which suggests a potential trade-off between generalization and specific performance, purely training in the simulation with inaccurate state occupancy distribution will limit the policy’s performance beforehand.
@@
Leveraging the data in the real world where the robot is directly situated is an intuitive solution. Nevertheless, it faces obstacles. Firstly, collecting real-world data is notoriously expensive while RL methods are extremely data-hungry. An applicable policy generally requires several months of real-world experience which is hardly affordable. Secondly, absent privileged knowledge (heightfield, contact force, mass, friction, etc) in real-world settings complicates tasks on tricky terrains. One example is the stairs, where a blind robot with only proprioceptive observations (i.e. joint position & velocity) needs extensive exploration to learn to lift its legs, while the one with precise height information can better find such a solution. On top of that, noisy observation also leads to training instability.
@@
Previous works have attempted to cope with these impediments by reshaping reward functions [11, 12], taking advantage of sample-efficient algorithms [11, 12, 13], or harnessing model-based RL [14]. However, these methods, trained directly in the real world, fail to produce superior locomotion policies and are vulnerable during training compared to zero-shot methods.
@@
Drawing inspiration from how animals use familiar environments to navigate new ones, we investigate to bridge the way back to the controllable simulation environment rather than trapping ourselves in the capricious and noisy real world. This approach offers all-around benefits. In one way, we can sidestep the problem of redesigning a delicate reward function corresponding to the limited observation, and take full advantage of a well-polished simulation training process [1, 3, 4, 15]. Additionally, it requires minimal real-world data while still living off the data abundance available in the simulation.
@@
In this work, we present LoopSR, an effective pipeline for lifelong policy adaptation in legged robots depicted in Fig. 1. At its core, LoopSR utilizes a transformer-based encoder [16] to leverage the latent space. By adopting an autoencoder architecture [17, 18] and contrastive loss [19, 20, 21, 22], the model extracts essential features for reconstructing simulation dynamics, ensuring both effectiveness and generalizability.
We implement learning-based and retrieval-based methods to derive simulation parameters from the latent variable, which compensate for each other in accuracy and robustness to out-of-distribution situations. Through comprehensive evaluations, LoopSR outperforms strong baselines in both the IsaacGym simulated environment [1] and the real-world scene.
@@
We proposed a novel lifelong policy adaptation method LoopSR that first incorporates real-world policy adaptation into simulated policy pretraining, achieving constant improvement during real-world deployments.
@@ II-A Representation Learning in RL
Representation learning [18, 23, 24, 25] has made up the bedrock for modern deep learning algorithms, and the field of RL and control is no exception. Previous works have made versatile attempts to take advantage of the representation learning and latent space [26, 27, 28], including contrastive learning [22, 29], reward function encoding [30, 31], domain/task-specific representation [17, 32, 33], and gait encodings [34, 35], boosting multi-modal input [24] or transferring through different tasks and embodiment [32].
@@ II-B Learning in the Real World
Continual learning in RL [38] is a daunting task as the result of the compounding error in drifted distribution. Works have been done to address the problem caused by protean data distribution through experience replay [39, 40], distillation [41], and reusing parts of model [42], etc. However, when it comes to real-world learning, the situation becomes progressively difficult due to noisy input, limited observation, and inaccessible reward functions.
[11, 12, 13] are among the few works tapping into this tricky problem. Instead of sim-to-real transfer, these works have made attempts to completely learn the policy in the real world by introducing reshaped reward functions and techniques of restricting action space to boost effective training. However, these algorithm needs careful designing in action space constraints and rewards to reluctantly obtain effective training and executable gaits, while failing to outstrip zero-shot methods.
With world models exhibiting competence in heterogeneous downstream tasks [43], [14] employ the Dreamer world model in the real-world learning task. The model-based method, though a practical alternative, still suffers from the problem of deficient privileged knowledge. The model’s accuracy in featuring real-world information is also worth questioning in comparison with the simulation, as model-based methods are susceptible to pervasive model biases. The vulnerability of training also places the risk that convergence is not guaranteed facing the composition of model biases and compounding error.
@@ III Methodology
As shown in Fig. 1, our proposed method LoopSR (Looping Sim and Real) supports lifelong policy adaptation in quadrupedal locomotion. In the following subsections, we first introduce the preliminaries (Sec. III-A) and provide an overview (Sec. III-B) of our algorithm. Then we illustrate our core training framework in Sec. III-C and provide important technical details in Sec. III-D. Finally, we present a script of theoretical analysis to support the validity (Sec. III-E).
III-A Preliminaries
We formulate the training process in the simulator as a Partially Observable Markov Decision Process (POMDP). We define the MDP in the simulation as , where denotes the state space, observation space, and action space respectively. denotes the transition function, and the reward function is denoted as . denotes the discount factor. We define the privileged environment parameters used to construct the simulation environment. Specifically, contains the knowledge of the terrain distribution (terrains are divided into 5 categories for feasibility) and robot-concerned parameters (mass, motor strength, etc.). The goal of the RL method is to get the policy to maximize the expected discounted return:
In the task of the lifelong policy adaptation, we need to adapt the pretrained policy to optimal in the real scene with the POMDP , where is implicit without access. In the real world, we can collect trajectories constantly but at an extremely low speed, at least times slower than in the simulation. In the later analysis, we take and to represent the training environment and the testing one.
@@ III-B Looping Sim and Real
Fig. 1 and pseudocode 1 outline LoopSR’s two-stage pipeline: pretraining and lifelong adaptation. During pretraining, LoopSR learns a deployable policy in the simulation while constructing an informative latent trajectory space by training an encoder-decoder network. In the adaptation stage, it encodes real-world trajectories into latent variables and refines simulation dynamics by leveraging stored rollouts and learned representations.
For pretraining, we employ Proximal Policy Optimization (PPO) [44] in the IsaacGym environment following [3]. The policy is based on DreamWaQ [45] and AMP [46, 47], utilizing an Asymmetric Actor-Critic architecture with expert motion data and domain randomizations [6, 7, 8] to enable real-world deployment. Meanwhile, LoopSR constructs a latent trajectory space by training the trajectory encoder (LoopSR-enc) alongside a multi-head decoder using collected rollouts and privileged information. These rollouts are stored as a reference dataset for later adaptation.
A collected real-world trajectory is encoded into a latent variable during the adaptation stage. The estimated is then estimated by combining retrieval from stored rollouts () and the decoded model output (). The final configures the IsaacGym simulation.
@@ III-C Network Architecture
Contrastive Loss. To guide the model to discriminate different terrains implicitly, we apply the contrastive loss in the form of supervised InfoNCE loss similar to [53]. The loss is defined to maximize the inner product of from the trajectories with the same terrain type label, whose theoretical background of maximizing the mutual information between positive samples and the latent representation is well-discussed in [20, 53].
where is the batch size and the superscript denotes the index of different trajectories.
Multi-head Decoder. We also design a multi-head decoder and to separately extract the terrain distribution and the robot proprioceptive parameters for disentanglement of different environment characteristics. We introduce , , and to represent the environment terrain distribution, robot parameter, and the randomization range of parameters in the training process. The losses are designed as follows:
where denotes the interval length of the randomization range. In Sec. IV-D, we conduct the ablation study to discuss the necessity of all loss terms.
@@ III-D Implementation Details
Soft Update. We implement the soft update of estimated environment parameters to avoid deleterious abrupt changes in terrain distribution and robot parameters. Specifically, the estimated environment parameters is updated through . is set to be . Such a method has the adaptive ability to attune to the environment setting which is not in the initial reference dataset. The necessity of soft update is ablated in Sec. IV-D.
Retrieval Process. Regarding training safety, we additionally employ the retrieval method to preclude drastically mistaken results in the face of out-of-distribution (o.o.d.) instances from learning-based methods. The final result is weighted as with set to be .
@@ III-E Theoretical Analysis
We provide a theoretical analysis here to explain the effectiveness of LoopSR. With the No Free Lunch Theorem, we demonstrate that wide domain randomization during policy training leads to suboptimal strategy in a specific environment. The influence on policy performance of the differences between the training environment and the testing one can be modeled as the discrepancy between optimal value functions. in the training environment and the testing one is derived in (5).
where
In (5), and are optimal value functions in and respectively, and as the optimal state value function on . informs the discrepancy in reward functions, which is negligible for the substantiated validity of from previous works [3, 4, 54, 15]. models the discrepancy caused by policy differences, which should approach with sufficient model capacity as and share the same input with detailed reasoning in [55].
The third term denotes the discrepancy induced by the shift in state occupancy distribution, where is closely related to the innate state occupancy distribution (e.g. the state distribution of the stair terrain differs from the plain one). However, wide domain randomization inherently leads to large with larger environment diversity. On the other hand, LoopSR introduces the process of building a digital twin of the real world and gradually training the policy on similar environments, thus efficiently mitigating such an error and optimizing the policy performance.
@@ IV Experiments
Terrain | Difficulty | Expert | Origin (DreamWaQ) | LoopSR (Ours) | RMA | SAC | World Model |
@@ IV-A Sim-to-Sim Experiment
Overview. Although simulation settings cannot perfectly replicate real-world complexities, they offer a controlled environment for conducting statistical experiments, facilitating comparisons when direct real-world computation is impractical. We leverage IsaacGym [1] for training and simulation, and the training code is developed based on [3].
Referred to Sec. III-B, the experiment consists of pertaining and adaptation stages. During pretraining, 4,096 robots are simulated across five distinct terrains (upward/downward slopes/stairs and plain), training in parallel for 30,000 iterations, with 24 simulation steps per iteration. During adaptation, data is collected in a testing environment chosen from three terrains, each with three difficulty levels (details see Table I), where the environment with the highest difficulty is an out-of-distribution one. Privileged knowledge is not accessed in the testing environment to better simulate real-world conditions.
In general, LoopSR runs sim-to-real loops for each testing environment. In each loop, we collect steps in the testing environment and train 500 iterations in the newly constructed training environment. In total, each policy is trained on steps of collected data, equivalent to half an hour in a real-world experiment. For a fair comparison, all baselines share the same total training iterations of 80,000. During evaluations, robots are commanded to walk forward at a velocity of to assess their maximum performance. The reward functions follow [3] and are averaged across simulation steps.
@@
Expert is a reference policy completely trained on the desired testing environment. The architecture follows DreamWaQ [45].
Origin (DreamWaQ) is a reference policy without refinement, representing a zero-shot sim-to-real workflow following DreamWaQ [45].
RMA is another zero-shot sim-to-real workflow following RMA [15]. Here, the teacher and student policy are trained for 80,000 iterations respectively.
SAC follows [11] with SAC used in pretraining and continual training.
Results. The simulation results are shown in Tab. I, where LoopSR achieves prominent enhancement and closely matches the performance of Expert, especially in challenging terrains like Stairs and Discrete. This outperformance over zero-shot sim-to-real methods (Origin, RMA) confirms LoopSR’s success in continuous optimization. At the same time, direct real-world training approaches (SAC, World Model) mainly fail, typically owning to their fragility to the shift in state occupancy distribution. The reward curve concerning the sim-to-real loops is shown in Fig. 3, where each sim-to-real loop integrates simulated training of iterations. We find out that LoopSR surpasses the performance bottleneck which troubles the original sim-to-real baseline, demonstrating the necessity of policy adaptation.